OpenStudy (anonymous):

Express the complex number in trigonometric form. -3i

OpenStudy (jdoe0001):

\(-3i\implies \begin{array}{ccllll} 0&-3i\\ \uparrow &\uparrow \\ x&y \end{array}\qquad \begin{cases} r=\sqrt{x^2+y^2}\\ \theta=tan^{-1}\left( \frac{y}{x}\right) \end{cases}\qquad 0-3i\implies r[cos(\theta)+i\ sin(\theta)]\) notice your "x" notice your "y" can you plot them both here ?

OpenStudy (jdoe0001):

by plotting them, you can get the angle easy, rather than getting the tangent, which in this case will be undefined anyway

OpenStudy (anonymous):

@jdoe0001 so would it be 3(cos 0° + i sin 0°)

OpenStudy (jdoe0001):

well... what's your "r"? what's your \(\theta?\)